### v1

# WorldMem: Long-term Consistent World Simulation with Memory

Abstract
World simulation has gained increasing popularity due to its ability to model virtual environments and predict the consequences of actions. However, the limited temporal context window often leads to failures in maintaining long-term consistency, particularly in preserving 3D spatial consistency. In this work, we present WorldMem, a framework that enhances scene generation with a memory bank consisting of memory units that store memory frames and states (e.g., poses and timestamps). By employing a memory attention mechanism that effectively extracts relevant information from these memory frames based on their states, our method is capable of accurately reconstructing previously observed scenes, even under significant viewpoint or temporal gaps. Furthermore, by incorporating timestamps into the states, our framework not only models a static world but also captures its dynamic evolution over time, enabling both perception and interaction within the simulated world. Extensive experiments in both virtual and real scenarios validate the effectiveness of our approach.
Figure 1: WorldMem enables long-term consistent world simulation with an integrated memory mechanism. (a) Previous world simulation methods typically face the problem of inconsistent world due to limited temporal context window size. (b) WorldMem  empowers the agent to explore diverse and consistent worlds with an expansive action space, e.g., crafting environments by placing objects like pumpkin light or freely roaming around. Most importantly, after exploring for a while and glancing back, we find the objects we placed are still there, with the inspiring sight of the light melting surrounding snow, testifying to the passage of time. Red and green boxes indicate scenes that should be consistent. Project page at https://xizaoqu.github.io/worldmem.
1 Introduction
World simulation has gained significant attention for its ability to model environments and predict the outcomes of actions [2, 5, 1, 7, 27, 34]. Recent advances in video diffusion models have further propelled this field, enabling high-fidelity rollouts of potential future scenarios based on user actions, such as navigating through an environment or interacting with objects. These capabilities make world simulators particularly promising for applications in autonomous navigation [7, 2] and as an alternative to traditional game engines [5, 27].
Despite impressive progress, existing methods suffer from a key limitation: they struggle to maintain long-term consistency. A prominent issue is the lack of 3D spatial consistency: As illustrated in Figure 1(a), when the viewpoint moves in one direction and then returns, the content of the environment has changed. This inconsistency arises because video diffusion models operate within limited temporal windows, restricting their ability to retain information over extended time periods. Ideally, the model should be able to recall the generated environment and past events even across long temporal gaps. However, simply increasing the temporal window is impractical due to the prohibitive memory and computational costs.
To address this limitation, we introduce WorldMem, a novel framework that enhances long-term consistency in video-based world simulators through a dedicated memory mechanism. Our core idea is to continuously store visual and state information in an external memory bank and retrieve the most relevant memories for generating new frames. By removing the constraint of a bounded time horizon, this memory mechanism allows the model to capture and reuse historical cues, ensuring better preservation of environment details and past events.
WorldMem is built upon a Conditional Diffusion Transformer (CDiT) [28] that integrates external action signals to guide first-person viewpoint generation. By leveraging Diffusion Forcing (DF) [3] during training, the model supports autoregressive generation, enabling extended temporal simulation. Unlike previous methods that store past content either as explicit 3D reconstructions [24, 36] or within implicit network weights [1, 17], we instead maintain a memory bank composed of memory units – each containing a past frame and its corresponding state. This approach offers both flexibility and precision: We avoid rigid 3D reconstructions while retaining the fine-grained visual details necessary for accurate scene synthesis. The memory bank is continuously updated throughout generation, and a dedicated retrieval strategy effectively retrieves relevant units to guide future generation steps.
To effectively incorporate the retrieved memory units, we adopt the flexible DF paradigm, treating memory frames as “clear” latents alongside noisy latents in the denoising loop. This design obviates the need for additional conditioning techniques [15, 48, 43] and prevents information loss that may occur when passing through separate conditioning modules. Following [8, 38], we model information exchange between memory frames and the frames currently being generated through attention [35], where the current frames act as queries and the memory frames serve as keys and values. To ensure robust information retrieval, we represent states with specialized embeddings that enrich both queries and keys. Specifically, we employ Plücker embeddings [31] for dense pose representation and incorporate relative embeddings to facilitate efficient learning.
We evaluate WorldMem on a customized Minecraft benchmark [6] and on RealEstate10K [51]. The Minecraft benchmark includes diverse terrains (e.g., plains, savannas, and deserts) and various action modalities (movement, viewpoint control, and event triggers). Extensive experiments show that WorldMem significantly improves 3D spatial consistency, enabling robust viewpoint reasoning and high-fidelity scene reconstructions, as shown in Figure 1(b). Furthermore, in dynamic environments, WorldMem accurately tracks and follows evolving events and environment changes, demonstrating its ability to both perceive and interact with the generated world.
Overall, WorldMem’s enhanced capacity to maintain long-term consistency represents a critical advancement for interactive world simulators, paving the way for more accurate, persistent, and immersive virtual environments. We hope our promising results will inspire future research on memory-based consistent world simulation.
2 Related Work
Video diffusion model. With the rapid advancement of diffusion models [33, 28, 3], video generation has made significant strides [37, 39, 4, 9, 26, 20, 44]. The field has evolved from traditional U-Net-based architectures [37, 4, 9] to Transformer-based frameworks [26, 25, 50], enabling video diffusion models to generate highly realistic and temporally coherent videos. Recently, autoregressive video generation [3, 21, 15] has emerged as a promising approach to extend video length, theoretically indefinitely. Notably, Diffusion Forcing [3] introduces a per-frame noise-level denoising paradigm. Unlike the full-sequence paradigm, which applies a uniform noise level across all frames, per-frame noise-level denoising offers a more flexible approach, enabling autoregressive generation.
Interactive world simulation. World simulation aims to model an environment by predicting the next state given the current state and action. This concept has been extensively explored in the construction of world models [11] for agent learning [10, 12, 13, 18]. With advances in video generation, high-quality world simulation with robust control has become feasible, leading to numerous works focusing on interactive world simulation [2, 5, 1, 7, 27, 34, 47]. These approaches enable agents to navigate generated environments and interact with them based on external commands.
However, due to context window limitations, such methods tend to forget previously generated content, leading to inconsistencies in the simulated world, particularly in maintaining 3D spatial coherence.
Consistent world simulation. Ensuring the consistency of a generated world is crucial for effective world simulation. Existing approaches can be broadly categorized into two types: explicit reconstructions and implicit conditioning.
One mainstream approach to ensuring consistency is to explicitly reconstruct the generated world into a 3D representation [24, 8, 36, 29, 46, 45, 23]. While this strategy can reliably maintain consistency, it imposes strict constraints on flexibility: Once the world is reconstructed, modifying or interacting with it becomes challenging.
Another research direction focuses on implicit learning. Methods like [1, 34] ensure consistency by overfitting to predefined scenarios (e.g., specific CS:GO or DOOM maps), limiting scalability. StreamingT2V [15] maintains long-term consistency by continuing on both global and local visual contexts from previous frames, while SlowFastGen [17] progressively trains LoRA [19] modules for memory recall. However, these methods rely on abstract representations, making accurate scene reconstruction challenging. In contrast, our approach retrieves information from previously generated frames and their states, ensuring world consistency without overfitting to specific scenarios.
3 WorldMem
This section details the methodology of WorldMem. Sec. 3.1 introduces the relevant preliminaries, while Sec. 3.2 describes the interactive world simulator serving as our baseline. Sec. 3.3 presents the core of our proposed memory mechanism, followed by Sec. 3.4 covering the memory retrieval strategies and Sec. 3.5 discussing the key designs for state embeddings in memory blocks.
Figure 2: Comprehensive overview of WorldMem. The framework comprises a conditional diffusion transformer integrated with memory blocks, with a dedicated memory bank storing memory units from previously generated content. By retrieving these memory units from the memory bank and incorporating the information by memory blocks to guide generation, our approach ensures long-term consistency in world simulation.
3.1 Preliminary
Video diffusion models. Video diffusion models generate video sequences by iteratively denoising Gaussian noise through a learned reverse process:
$p_{\theta}(\mathbf{x}^{k-1}_{t}|\mathbf{x}^{k}_{t})=\mathcal{N}(\mathbf{x}^{k-1}_{t};\mu_{\theta}(\mathbf{x}^{k}_{t},k),\sigma_{k}^{2}\mathbf{I}),$ (1)
where all frames $(\mathbf{x}_{t}^{k})_{1\leq t\leq T}$ share the same noise level $k$, and $T$ is the context window length. This full-sequence approach enables global guidance but lacks flexibility in sequence length and autoregressive generation.
Autoregressive video generation. Autoregressive video generation aims to extend videos over the long term by predicting frames sequentially [22, 41]. While various methods exist for autoregressive generation, Diffusion Forcing (DF) [3] provides a neat and effective approach to achieve this. Specifically, DF introduces per-frame noise levels $k_{t}$:
$p_{\theta}(\mathbf{x}^{k_{t}-1}_{t}|\mathbf{x}^{k_{t}}_{t})=\mathcal{N}(\mathbf{x}^{k_{t}-1}_{t};\mu_{\theta}(\mathbf{x}^{k_{t}}_{t},k_{t}),\sigma_{k_{t}}^{2}\mathbf{I}),$ (2)
Unlike full-sequence diffusion, DF generates video flexibly and stably beyond the training horizon. Autoregressive generation is a special case when only the last one or a few frames are noisy. With autoregressive video generation, long-term interactive world simulation becomes feasible.
3.2 Interactive World Simulation
Before introducing the memory mechanism, we first outline the construction of a world simulator that incorporates actions as additional inputs within a conditional diffusion transformer framework. Following previous works [5], we adopt a conditional diffusion transformer as the baseline for interactive world simulation. Specifically, we adopt a DiT [28] architecture for video generation and use DF [3] for autoregressive generation. As illustrated in Figure 2 (a), our model comprises multiple DiT blocks with spatial and temporal modules for spatiotemporal reasoning. In the temporal module, causal attention ensures each frame only attends to preceding frames.
To further support interactive world simulation, we integrate interactive modules that incorporate external control signals into the models. The primary control signal is action [27, 5, 47], which is crucial in guiding the simulation. In our Minecraft experiments, the action space consists of $25$ dimensions, including movements (left, right, forward, back), view controls (look up, look down), and event triggers (object placement, item switching). To condition the model on these actions, we employ a multi-layer perceptron (MLP) to project the control signals into the embedding space. As shown in Figure 2 (c), we add the action embeddings and denoising timestep embedding together and inject the information to temporal blocks by AdaLN  [42], following the paradigm of previous approaches [2, 5]. Please note that we also inject timestep embeddings to spatial blocks in the same way, but we do not highlight that in Figure 2 for simplicity. Likewise, default attention settings like residual connections, multi-head attention, and feedforward networks are also omitted in Figure 2 for simplicity.
With the combination of conditional DiT [28] and DF [3] as a baseline, we can effectively simulate a long-term interactive world. However, due to the high computational cost of video generation, the context window remains limited. Consequently, content outside this window is forgotten, leading to inconsistencies in long-term generation [5].
3.3 Condition on Memory
We introduce a memory mechanism to prevent the model from forgetting content outside the context window. Intuitively, this mechanism functions as a persistent external buffer, allowing the model to revisit past frames and maintain a continuous record of the evolving scene. Specifically, we design a memory bank that continuously stores previously generated results. The memory is structured as a series of memory units $\mathbf{M}=\{\mathbf{I},\mathbf{p},\mathbf{t}\}$, which include visual content in the form of memory frames $\mathbf{I}$, along with state information such as poses $\mathbf{p}$ and timestamps $\mathbf{t}$.
However, effectively utilizing these memory units to guide generation is non-trivial. Unlike existing conditioning methods [15, 17], which condition on high-level and abstract features, our objective is to accurately reconstruct previously seen scenarios, even under significant viewpoint changes or scene changes. This requires the conditioning module to reason about the relationships between past results and the current generation.
To achieve this, we adopt the cross-attention mechanism [35] to guide information exchange between the denoising frames and the memory frames. Let $\mathbf{X}\in\mathbb{R}^{l_{1}\times d}$ be the denoising frames (queries) and $\mathbf{M}\in\mathbb{R}^{l_{2}\times d}$ be the memory frames (keys and values), where $l_{1}$ and $l_{2}$ are the flattened lengths of $\mathbf{X}$ and $\mathbf{M}$, respectively, and $d$ is the channel dimension. We first incorporate the state embeddings $\mathbf{E}_{X}$ and $\mathbf{E}_{M}$ by summation:
$\mathbf{\tilde{X}}=\mathbf{X}+\mathbf{E}_{X},\quad\mathbf{\tilde{M}}=\mathbf{M}+\mathbf{E}_{M}.$ (3)
We then compute cross-attention as:
$\mathbf{X}^{\prime}=\text{CrossAttn}(\text{Q}=p_{q}(\mathbf{\tilde{X}}),\text{K}=p_{k}(\mathbf{\tilde{M}}),\text{V}=p_{v}(\mathbf{M})),$ (4)
where $\mathbf{X}^{\prime}$ is the output after attending to relevant information in the memory frames, $p_{q}$, $p_{k}$, and $p_{v}$ are attention projectors. This allows the network to focus on memory content that is most pertinent to the current generation step. We will further elaborate on how $\mathbf{E}_{X}$ and $\mathbf{E}_{M}$ are computed in Sec. 3.5.
To incorporate memory units into different DiT blocks, we leverage the advantages of DF by treating memory frames as clear frames and feeding them into the diffusion pipeline alongside input noisy frames, as shown in Figure 2 (b). For the training process, all frames in the temporal context window have different levels of noise, and these memory frames have a clear level of noise (indicated as $k_{min}$):
$\mathbf{x}_{i}^{k},k\begin{cases}\sim\mathcal{P}(\mathbf{k}),&i\in\mathbf{TF}\\ =k_{\text{min}},&i\in\mathbf{MF}\end{cases}$ (5)
where $\mathbf{x}_{i}^{k}$ indicates the $i^{\text{th}}$ frame with noise level $k$, $\mathbf{k}$ is the set of possible noise levels, $\mathbf{TF}$ is the set of frame indices within temporal context window, $\mathbf{MF}$ is the set of frame indices in the memory frame part. For inference,
$\mathbf{x}_{i}^{k},k=\begin{cases}k_{\text{max}},&i\in\mathbf{GF}\\ k_{\text{min}},&\text{otherwise}\end{cases}$ (6)
where $\mathbf{GF}$ is the set of indices of generated frames, $k_{\text{max}}$ indicates the max level of noise.
We ensure that these memory frames influence only the memory block by masking them in the temporal block. As shown in Figure 2 (c), given the memory frame length as $L_{M}$, the mask can be noted as:
$A_{\text{mask}}(i,j)=\begin{cases}1,&\text{if }i\leq L_{M}\text{ and }j=i\\ 1,&\text{if }i>L_{M}\text{ and }j\leq i\\ 0,&\text{otherwise}\end{cases}$ (7)
Within the memory block, we use cross-attention to retrieve information from memory frames and guide the generation of the current frame.
3.4 Memory Retrieve
Since the number of memory frames available for conditioning is limited, an efficient strategy is required to sample memory units from the memory bank. We adopt a greedy matching algorithm based on frame-pair similarity, where similarity is defined using the field-of-view (FOV) overlap ratio and timestamp differences as confidence measures. Algorithm 1 presents our approach to memory retrieval. Additionally, we apply a similarity filter to eliminate redundant units, ensuring that only the most relevant frames are retrieved for reference in the current generation. Figure 3 visualizes the overlapping metric with the Monte Carlo sampling.
Input:
  • A memory bank of $n$ historical states $\{(\mathbf{I}_{i},\mathbf{p}_{i},t_{i})\}_{i=1}^{n}$.
  • Current state $(\mathbf{I}_{c},\mathbf{p}_{c},t_{c})$.
  • Memory condition length $L$.
  • Similarity threshold $tr$.
  • Weights $w_{o},w_{t}$.
Output: A list of selected state indices $S$.
1
2begin
3       $\Delta$ Compute Confidence Score
4       Compute FOV overlap ratio $\mathbf{o}$ with the current state via Monte Carlo sampling.
5       Compute time difference $\mathbf{d}=\text{Concat}(\{|t_{i}-t_{c}|)\}_{i=1}^{n}$.
6       Compute confidence $\boldsymbol{\alpha}=\mathbf{o}\cdot w_{o}$ $\mathbf{d}\cdot w_{t}$.
7
8      $\Delta$ Selection with Similarity Filtering
9       Initialize $S=\varnothing$.
10       for $m=1$ to $L$ do
11             Select the state $i^{*}$ with the highest $\alpha_{i^{*}}$.
12             Append $i^{*}$ to $S$.
13             Remove all states $j$ whose feature similarity with $i^{*}$ exceeds $tr$.
14
15       end for
16      Return $S$.
17 end
18
Algorithm 1 Memory Retrieval Algorithm
Figure 3: Two-view FOV overlapping visualization
3.5 State Embedding Design
The design of state embedding is critical for the memory block to retrieve valid information effectively.
A key aspect of this process is capturing spatial information, which primarily depends on pose embeddings. Inspired by [14, 8], which transforms frame-wise poses into dense positional embeddings to encode fine-grained spatial details, we adopt Plücker embedding [30] as our pose representation. For timestamp embeddings, a simple MLP-based mapping is sufficient. The state embedding $\mathbf{E}$ is then computed as the sum of pose and timestamp embeddings:
$\mathbf{E}=G_{p}(\text{PE}(\mathbf{p}))+G_{t}(t),$ (8)
where $G_{p}$ and $G_{t}$ are MLP layers mapping vectors to embeddings, and PE denotes the Plücker embedding function, which maps $\mathbf{p}\in\mathbb{R}^{5}$ (x, y, z, pitch, yaw) to $\mathbf{p}\in\mathbb{R}^{h\times w\times 6}$.
Additionally, we observe that relative embeddings are more effective than absolute embeddings for spatial reasoning, as they significantly reduce the model’s learning complexity. To leverage this, we introduce relative embedding in the memory attention block, where the Query embedding is always set to zero, and the Key embedding is derived from the relative pose. To integrate relative embeddings into our architecture, we separate the cross-attention into frame-wise querying, where input frames attend to memory frames independently. This design is illustrated in Figure 2 (d).
4 Experiments
Figure 4: Qualitative results. We showcase WorldMem’s capabilities through two sets of examples. Top: A comparison with Ground Truth (GT). WorldMem accurately models diverse dynamics (e.g., rain) by conditioning on 600 past frames, ensuring temporal consistency. Bottom: Interaction with the world. Objects like hay in the desert or wheat in the plains persist over time, with wheat visibly growing. For the best experience, see the supplementary videos.
Datasets. We use MineDojo [6] to create diverse training and evaluation datasets in Minecraft, configuring diverse environments (e.g., plains, savannas, ice plains, and deserts), agent actions, and interactions. For real-world scenes, we utilize RealEstate10K [51] with camera pose annotations, designing trajectories that incorporate past scenes to evaluate long-term world consistency.
Metrics. For quantitative evaluation, we employ reconstruction metrics, where the method of obtaining ground truth (GT) varies by specific settings. We then assess the consistency and quality of the generated videos using PSNR, LPIPS [49], and reconstruction FID (rFID) [16], which collectively measure pixel-level fidelity, perceptual similarity, and overall realism.
4.1 Results on Generation Benchmark
Comparisons on Minecraft Benchmark. We compare our approach with a standard full-sequence training method [14, 40] and Diffusion Forcing (DF) [3]. The key differences are as follows: the full-sequence conditional diffusion transformer [28] maintains the same noise level during training and inference, DF introduces different noise levels for training and inference, and our method incorporates a memory mechanism. To assess both short-term and long-term world consistency, we conduct evaluations within and beyond the context window. We evaluate both settings on 300 test videos and use 4,800 frames to compute rFID. In the following experiments, the agent’s poses are generated by the game simulator as ground truth. However, in real-world scenarios, only the action input is available, and the pose is not directly observable. In such cases, the next-frame pose can be predicted based on the previous scenes, past states, and the upcoming action. We explore this design choice in the supplementary material.
Within context window. For this experiment, all methods use a context window of 16, while our approach additionally maintains a memory window of 8. We test on customized motion scenarios (e.g., turn left, then turn right or move forward, then backward) to assess self-contained consistency, where the ground truth consists of previously generated frames at the same positions. As shown in Table 1 and Figure 5, the full-sequence baseline suffers from inconsistencies even within its own context window. DF improves consistency by enabling greater information exchange among generated frames. Our memory-based approach achieves the best performance, demonstrating the effectiveness of integrating a dedicated memory mechanism.
Beyond context window. In this setting, all methods use a context window of 8 and generate 100 future frames; our method further employs a memory window of 8 while initializing a 600-frame memory bank. We compute the reconstruction error using the subsequent 100 ground truth frames after 600 frames. Full-sequence methods can not roll out that long so we exclude it. DF exhibits poor PSNR and LPIPS scores, indicating severe inconsistency with the ground truth beyond the context window. Additionally, its low rFID suggests notable quality degradation. In contrast, our memory-augmented approach consistently outperforms others across all metrics, demonstrating superior long-term consistency and quality preservation. Figure 6 further substantiates these findings.
Figure 5: Within context window evaluation examples. It illustrates an example where the motion sequence first involves turning right and then returning to the original position, demonstrating methods’ ability to maintain self-contained consistency.
Table 1: Evaluation on Minecraft benchmark [6]
Within context window
Methods PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Full Sequence 20.35 0.0691 13.87
Diffusion Forcing [3] 26.56 0.0094 13.88
Ours 27.01 0.0072 13.73
Beyond context window
Methods PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Full Sequence / / /
Diffusion Forcing [3] 18.04 0.4376 51.28
Ours 25.32 0.1429 15.37
Figure 6: Beyond context window evaluation examples. It shows that Diffusion-Forcing suffers from inconsistency and quality degradation after generating a certain number of frames. In contrast, our method maintains high quality and faithfully reconstructs previously observed scenarios.
Comparisons on Real Scenarios. We further compare our approach with DFoT [32] on the RealEstate dataset [51]. Similar to DF [3], DFoT discards previously generated content that lies outside its context window and fails to maintain global world consistency.
To verify this, we currate 100 samples and test the 360-degree consistency of both DFoT and our method by rotating the scene a full 360 degrees and checking whether it returns to the same position. We measure the quality of reconstruction by comparing the first frame and the final frame (after a full rotation). As shown in Table 2, our method outperforms DFoT in all metrics, indicating that our integrated memory mechanism substantially improves consistency. Qualitative comparisons in Figure 7 further support these findings.
Figure 7: Examples on RealEstate [51]. DFoT [32] discards content beyond its context window, losing 360-degree consistency. In contrast, our method preserves details and accurately returns to the original location.
Table 2: Evaluation on RealEstate10K [51]
Methods PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
DFoT [32] 8.396 0.6676 156.74
Ours 20.19 0.1773 67.14
Qualitative results. Figure 4 showcases WorldMem’s capabilities. The top section demonstrates its ability to operate in a free action space across diverse environments. Given a 600-frame memory bank, our model generates 100 future frames while preserving the ground truth’s actions and poses, ensuring strong world consistency. The bottom section highlights dynamic environment interaction. By using timestamps as embeddings, the model remembers environmental changes and captures natural event evolution, such as plant growth over time.
4.2 Ablation
Figure 8: Long-term Generation Comparison. This figure presents the PSNR of different ablation methods compared to the ground truth over a 300-frame sequence. The results show that our method without memory blocks or using random memory retrieval exhibits immediate inconsistencies with the ground truth. Additionally, the model lacking relative embeddings begins to degrade significantly beyond 100 frames. In contrast, our full method maintains strong consistency even beyond 300 frames.
Embedding designs. The design of embeddings within the memory block is crucial for effective cross-frame relationship modeling. We evaluate three strategies (Table 3): (1) sparse pose embedding with absolute encoding, (2) dense pose embedding with absolute encoding, and (3) dense pose embedding with relative encoding. Results show that dense pose embeddings (Plücker embedding) significantly enhance all metrics, emphasizing the benefits of richer pose representations. Switching from absolute to relative encoding further improves performance, particularly in LPIPS and rFID, by facilitating relationship reasoning and information retrieval. As illustrated in Figure 8, absolute embeddings accumulate errors over time, while relative embeddings maintain stability even beyond 300 frames.
Table 3: Ablation on embedding designs
Pose type Embed. type PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Sparse Absolute 20.67 0.2887 39.23
Dense Absolute 23.63 0.1830 29.34
Dense Relative 25.32 0.1429 15.37
Figure 9: Results w/o and w/ time condition. Without timestamps, the model fails to differentiate memory units from the same location at different times, causing errors. With time conditioning, it aligns with the updated world state, ensuring consistency.
Time condition. We ablate the effectiveness of timestamp embedding in Table 4. For this experiment, we curate 100 video samples featuring placing events and evaluate whether future generations align with event progression. As shown in the table, incorporating time embeddings significantly improves PSNR and LPIPS, indicating that adding temporal information helps the model faithfully reproduce event changes in world simulation. Since events like plant growth are inherently unpredictable, we do not conduct quantitative evaluations on such cases but instead provide qualitative illustrations in Figure 9.
Table 4: Ablation on time condition
Time embedding PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
w/o 23.17 0.1989 23.89
w/ 25.12 0.1613 16.53
Memory retrieve strategy. We analyze memory retrieval strategies in Table 5. Random sampling from the memory bank leads to poor performance and severe quality degradation, as evidenced by a sharp drop in rFID and rapid divergence from the ground truth (Figure 8). The confidence-based filtering significantly enhances consistency and generation quality. Additionally, we refine retrieval by filtering out redundant memory units based on similarity, further improving all evaluation metrics and demonstrating the effectiveness of our approach.
Table 5: Ablation on memory retrieve strategy
Strategy PSNR $\uparrow$ LPIPS $\downarrow$ rFID $\downarrow$
Random 18.32 0.3224 47.35
+ Confidence Filter 23.12 0.1863 24.33
+ Similarity Filter 25.32 0.1429 15.37
5 Conclusion
In conclusion, WorldMem tackles the longstanding challenge of maintaining long-term consistency in world simulation by employing a memory bank of past frames and associated states. Its memory attention mechanism enables accurate reconstruction of previously observed scenes – even under large viewpoints or temporal gaps – and effectively models dynamic changes over time. Extensive experiments in both virtual and real settings confirm WorldMem’s capacity for robust, immersive world simulation. We hope our work will encourage further research on the design and applications of memory-based world simulators.